Perform Calculations Using Charless Law CONTENT FEEDBACK Question If a balloon at 25C with 2.3 L of volume is expanded to 40.0 L, what will the new Celsius temperature be? Report your answer with the correct number of significant figures. Provide your answer below: degrees C
Answer #1

Using charles Law

\frac{V_1}{T_1} = \frac{V_2}{T_2}

V1 = 2.3L

T1 = 273 + 25 = 298K

V2 = 40.0L

T_2 = \frac{V_2T_1}{V_1} = \frac{40.0 * 298}{2.3} = 5182.60K

Temperature in Celsius Scale = 5182.60 - 273 = 4909.6

Now since V1 and T1 is having volume and temperature till 2 significant digits, hence the final answer must be with 2 signficant digits,

Hence the answer will be 4900C, which is 4.9 * 10^3 C